# From Limited Labels to Open Domains:An Efficient Learning Method for Drone-view Geo-Localization

### v2

@@ Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, the CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot and cross-domain initialization.
Index Terms:
drone-view geo-localization, structural invariance learning, spatial invariance learning, knowledge transfer.
@@ I Introduction
In this paper, we propose an efficient learning network with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by spatial ambiguity, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
@@
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed method is presented in detail. The experimental results are reported and analyzed in Section LABEL:result. Finally, conclusions are outlined in Section LABEL:conclusion.
@@ II-B Unsupervised Cross-Domain Retrieval
In recent years, cross-domain image retrieval based on the supervised methods has made significant progress [10314802, 10812860, 9963608, 10855601]. However, the reliance on large-scale annotated data has limited its applicability, prompting researchers to explore unsupervised cross-domain retrieval methods [10109197, 9783116, 9891821]. Current methods [Chen_2021_ICCV, Xuan_2021_CVPR, dai2022cluster] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [ester1996density] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [yang2022augmented, yang2024shallow] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
To better learn cross-view correlations, we adopt a pseudo-label-based training scheme using unpaired data, where pseudo-labels are generated through clustering, similar to the methods discussed previously. In contrast, we initialize the model with a small amount of paired data. It leads to a more compact and discriminative feature space and enables more reliable pseudo-label generation. Notably, although unsupervised DVGL can also be considered an unsupervised cross-domain retrieval task, existing methods remain limited and have been discussed in the introduction. We refrain from repeating the details here.
@@ III Proposed Method
Problem Formulation: Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone image $x_{i}^{d}$ is paired with a corresponding satellite image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
@@ III-A1 Structural Invariant Learning (SIFL) Block
Subsequently, a depthwise separable $3\times 3$ convolution $\varphi^{3}_{q}$ is used to mine local structural priors, while residual connections preserve the original topological information. After normalization $\mathcal{N}(\cdot)$, the feature representation $\mathbf{f}_{v+}^{u}$ can be obtained. To further enhance the capability of structural invariance feature learning, we proposed a channel-wise refinement (CWR) module. Specifically, we project $\mathbf{f}_{v+}^{u}$ into the latent space through two cascaded linear transformation layers $\mathcal{I}(\cdot)$ to learn high-order interactions between channels. A subsequent $1\times 1$ convolution $\psi^{1}_{q}$ is applied to explicitly capture local structural features.
@@
$\mathbf{f}_{v+}^{u++}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+})$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}(\mathcal{I}^{+}(\mathcal{A}(\mathbf{f}_{v++}^{u+})))\times\mathcal{P}(\mathbf{f}_{v+}^{u++})+\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+}.$ (4)
@@ III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
In the spatial invariance enhancement learning module, we preserve the dimensions, namely height $H$ and weight $W$ of the input features. By permuting the input features $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$ along the channel-height (C-H) and channel-width (C-W) dimensions, we obtain $\mathbf{f}_{v}^{hu}\in\mathbb{R}^{H\times C\times W}$ and $\mathbf{f}_{v}^{wu}\in\mathbb{R}^{W\times H\times C}$, respectively. This allows the module to capture multi-perspective dependencies from different dimensions. Subsequently, global feature representations $\mathbf{f}_{v2}^{hu}\in\mathbb{R}^{2\times H\times W}$ and $\mathbf{f}_{v2}^{wu}\in\mathbb{R}^{2\times H\times W}$ are generated by concatenating the results of max pooling and average pooling applied to the permuted features. A $3\times 3$ depthwise separable convolution $\varphi^{3}_{v}$ is then employed to obtain a more compact feature representation. Subsequently, a strong cross-interaction across the C-H and C-W is performed, followed by separate $1\times 1$ convolution $\psi^{1}_{k}$ to further enhance the spatial invariance of the feature representation. Finally, the fused features are normalized using a sigmoid function $\mathcal{F}$ to generate attention weights, which are multiplied with the original feature maps and restored to the original dimensions $\mathbb{R}^{C\times H\times W}$, yielding the spatially enhanced feature $\mathbf{f}_{v+}^{hu}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{f}_{v+}^{wu}\in\mathbb{R}^{C\times H\times W}$ as
$\mathbf{f}_{v+}^{hu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu}))+\psi^{1}_{k}(\mathbf{f}_{v2}^{wu}))\times\mathbf{f}_{v}^{u}$ (5)
$\mathbf{f}_{v+}^{wu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu}))+\psi^{1}_{k}(\mathbf{f}_{v2}^{hu}))\times\mathbf{f}_{v}^{u}.$ (6)
@@ III-A3 Multi-level Optimization
[…]
@@ III-B2 Cross-View Feature Memory Updating
$\quad\phi_{k}^{s(\omega)}\leftarrow\alpha\phi_{k}^{s(\omega-1)}+(1-\alpha)q_{s}$ (10)

### v3

@@ Abstract
Traditional supervised drone-view geo-localization (DVGL) methods heavily depend on paired training data and encounter difficulties in learning cross-view correlations from unpaired data. Moreover, when deployed in a new domain, these methods require obtaining the new paired data and subsequent retraining for model adaptation, which significantly increases computational overhead. Existing unsupervised methods have enabled to generate pseudo-labels based on cross-view similarity to infer the pairing relationships. However, geographical similarity and spatial continuity often cause visually analogous features at different geographical locations. The feature confusion compromises the reliability of pseudo-label generation, where incorrect pseudo-labels often drive negative optimization. Given these challenges inherent in both supervised and unsupervised DVGL methods, we propose a novel cross-domain invariant knowledge transfer network (CDIKTNet) with limited supervision, whose architecture consists of a cross-domain invariance sub-network (CDIS) and a cross-domain transfer sub-network (CDTS). This architecture facilitates a closed-loop framework for invariance feature learning and knowledge transfer. The CDIS is designed to learn cross-view structural and spatial invariance from a small amount of paired data that serves as prior knowledge. It endows the shared feature space of unpaired data with implicit cross-view correlations at initialization, which alleviates feature confusion. Based on this, the CDTS employs dual-path contrastive learning to further optimize each subspace while preserving consistency in a shared feature space. Extensive experiments demonstrate that CDIKTNet achieves state-of-the-art performance under full supervision compared with those supervised methods, and further surpasses existing unsupervised methods in both few-shot initialization and cross-domain initialization setting.
@@ I Introduction
In this paper, we propose an efficient learning method with limited supervision for DVGL that mitigates the adverse effects of inaccurate pseudo-labels caused by feature confusion, called the cross-domain invariant knowledge transfer network (CDIKTNet). Specifically, the CDIKTNet consists of two core components, including the cross-domain invariance sub-network (CDIS) and the cross-domain transfer sub-network (CDTS). The CDIS utilizes a small amount of paired data to learn latent cross-view structural and spatial invariance, which serves as prior knowledge. As illustrated in Fig. 2(b), it endows the shared feature space of unpaired data with similar implicit cross-view correlations at initialization to alleviate feature confusion. Based on this, CDTS introduces dual-path contrastive learning to independently optimize subspaces under different views. Despite being optimized separately, these subspaces maintain alignment in the shared feature space due to the shared invariance prior. The main contributions can be summarized as follows.
@@
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
@@ II-B Unsupervised Cross-Domain Retrieval
In recent years, cross-domain image retrieval based on the supervised methods has made significant progress [9963608, 10897882]. CSCL [10812860] enhances cross-domain pedestrian matching by jointly leveraging structural alignment and semantic alignment to extract modality-invariant semantic cues. CSDN [10855601] mitigates the modality discrepancies by integrating visual features with high-level semantic information captured by CLIP [radford2021learning] to enrich the cross-domain representation space. To address the severe performance degradation caused by occlusion, OAFR [10314802] simulates diverse occlusion patterns to enable the model to perceive OTP and PTP occlusions, and further reconstructs the occluded query features using the unoccluded gallery features, which leads to more robust cross-domain retrieval performance. However, the reliance on large-scale annotated data has limited its applicability. Fortunately, recent advances in domain adaptation [10962298, liu2025nonconvex, li2022cross] and semi/unsupervised learning [sun2021semi, jiang2022semi, du2021cross] have demonstrated that robust representations can be learned even under weak or missing labels. This research trend provides important insights for extending cross-domain retrieval toward unsupervised settings and has motivated researchers to explore unsupervised cross-domain retrieval methods [10109197, 9783116, 9891821]. Current methods [Chen_2021_ICCV, Xuan_2021_CVPR, dai2022cluster] utilized powerful feature backbone networks for feature initialization, employing clustering algorithms [ester1996density] to generate pseudo-labels and feature queues that provide stable negative samples for optimizing feature learning. However, these methods are primarily designed for homogeneous data with strong correlations, and their effectiveness tends to degrade significantly when applied to heterogeneous scenarios. To address the limitations, some researchers [yang2022augmented, yang2024shallow] have proposed cross-modal person retrieval methods based on dual-path contrastive learning. Nevertheless, these methods still depend on pre-trained person retrieval models for feature extraction during initialization.
To better learn cross-view correlations, we adopt a pseudo-label-based training scheme using unpaired data, where pseudo-labels are generated through clustering in accordance with the methods described above. In contrast, the model is initialized with a small amount of paired data. It leads to a more compact and discriminative feature space and enables more reliable pseudo-label generation. Notably, although unsupervised DVGL can also be considered an unsupervised cross-domain retrieval task, existing methods remain limited and have already been discussed in the Introduction. We refrain from repeating the details here.
@@ III Proposed Method
Problem Formulation. Given a drone-view image set $\mathcal{X}^{d}=\{x_{i}^{d}\}_{i=1}^{N}$ and a satellite-view image set $\mathcal{X}^{s}=\{x_{j}^{s}\}_{j=1}^{M}$, where each drone image $x_{i}^{d}$ is paired with a corresponding satellite image $x_{j}^{s}$ as a positive pair $(x_{i}^{d},x_{j}^{s})$, the objective is to learn a shared feature space $\mathcal{V}$ that minimizes the distance between positive pairs while maximizing the distance between negative pairs. Existing methods heavily rely on a large amount of paired data as positive samples for training to achieve high performance. In contrast, our method fundamentally differs by mitigating this dependency. The following sections provide a detailed explanation of our method.
@@ III-A1 Structural Invariant Learning (SIFL) Block
Subsequently, a depthwise separable $3\times 3$ convolution $\varphi^{3}_{q}$ is used to mine local structural priors, while residual connections preserve the original topological information. After normalization $\mathcal{N}(\cdot)$, the feature representation $\mathbf{f}_{v+}^{u}$ can be obtained. To further enhance the capability of structural invariance feature learning, we proposed a channel-wise refinement (CWR) module. Specifically, we project $\mathbf{f}_{v+}^{u}$ into the latent space through two cascaded linear transformation layers $\mathcal{I}(\cdot)$ to learn high-order interactions between channels. A subsequent $1\times 1$ convolution $\psi^{1}_{q}$ is applied to explicitly learn local structural features.
@@
$\mathbf{f}_{v++}^{u+}=\mathcal{N}(\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+})$ (3)
$\mathbf{f}_{v++}^{u++}\!\!=\!\!\mathcal{F}(\mathcal{I}^{+}(\mathcal{A}(\mathbf{f}_{v++}^{u+})))\times\mathcal{P}(\mathbf{f}_{v++}^{u+})+\varphi^{3}_{q}(\mathbf{f}_{v+}^{u+})+\mathbf{f}_{v+}^{u+}.$ (4)
@@ III-A2 Spatial Invariant Enhancement Learning (SIEL) Module
In the spatial invariance enhancement learning module, we preserve the dimensions, namely height $H$ and weight $W$ of the input features. By permuting the input features $\mathbf{f}_{v}^{u}\in\mathbb{R}^{C\times H\times W}$ along the channel-height (C-H) and channel-width (C-W) dimensions, we obtain $\mathbf{f}_{v}^{hu}\in\mathbb{R}^{H\times C\times W}$ and $\mathbf{f}_{v}^{wu}\in\mathbb{R}^{W\times H\times C}$, respectively. This allows the module to capture multi-perspective dependencies from different dimensions. Subsequently, global feature representations $\mathbf{f}_{v2}^{hu}\in\mathbb{R}^{2\times C\times H}$ and $\mathbf{f}_{v2}^{wu}\in\mathbb{R}^{2\times W\times C}$ are generated by concatenating the results of max pooling and average pooling applied to the permuted features. A $3\times 3$ depthwise separable convolution $\varphi^{3}_{v}$ and a separate $1\times 1$ convolution $\psi^{1}_{k}$ are then employed to obtain a more compact feature representation. Subsequently, cross-interaction across the C-H and C-W is performed to further enhance the spatial invariance of the feature representation. Finally, the fused features are normalized using a sigmoid function $\mathcal{F}$ to generate attention weights, which are multiplied with the original feature maps and restored to the original dimensions $\mathbb{R}^{C\times H\times W}$, yielding the spatially enhanced feature $\mathbf{f}_{v+}^{hu}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{f}_{v+}^{wu}\in\mathbb{R}^{C\times H\times W}$ as
$\mathbf{f}_{v+}^{hu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu}))+\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu})))\times\mathbf{f}_{v}^{hu}$ (5)
$\mathbf{f}_{v+}^{wu}=\mathcal{F}(\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{wu}))+\psi^{1}_{k}(\varphi^{3}_{v}(\mathbf{f}_{v2}^{hu})))\times\mathbf{f}_{v}^{wu}.$ (6)
@@ III-A3 Multi-level Optimization
In the CDIS, the features $\mathbf{f}_{v++}^{u++}$ obtained from the SIFL are optimized using the mean squared error loss $\mathcal{L}_{MSE}$ [xia2024enhancing], which improves the stability and consistency of the learned features. Furthermore, to enhance structural representation in the spatial invariance enhancement learning module, the features $\mathbf{f}_{v}^{au}$ are first extracted by the backbone and processed via global average pooling. The features $\mathbf{f}_{v}^{au}$ combined with $\mathbf{f}_{v+}^{hu}$ and $\mathbf{f}_{v+}^{wu}$, and the resulting combination are passed through a classifier module [shen2023mccg] to obtain $\mathbf{f}_{v+}^{cu}$. Subsequently, $\mathbf{f}_{v+}^{cu}$ are optimized with cross-entropy loss $\mathcal{L}_{CE}$ to enhance feature discrimination and cross-view consistency. In addition, to further enhance the feature representations, $\mathbf{f}_{v}^{au}$ undergo independent optimization using InfoNCE loss [oord2018representation] $\mathcal{L}_{InfoNCE}$. The final objective is to minimize the sum of all losses. It can be formulated as
@@ III-B1 Cross-View Feature Memory Initialization
Remark 2
In selecting the clustering algorithm, we refer to several unsupervised cross-domain retrieval works [dai2022cluster, adca] and adopt DBSCAN [ester1996density] for clustering. Unlike partition-based methods such as K-means [likas2003global], DBSCAN utilizes a density-based method, which handles clusters of arbitrary shapes and noisy data. It does not require a predefined number of clusters, nor does it force all points to be assigned to a cluster. Additionally, DBSCAN can identify and exclude noise points, which ensures the selection of valid samples even in cross-view challenges
@@ III-B2 Cross-View Feature Memory Updating
$\quad\phi_{l}^{s(\omega)}\leftarrow\alpha\phi_{l}^{s(\omega-1)}+(1-\alpha)q_{s}$ (10)